OpenStudy (anonymous):

College Algebra: reviewing for Final Exam!!!! Find the vertical asymptote(s):(given below)

OpenStudy (anonymous):

\[f(x)=\frac{ 4x-9 }{ 2x ^{2}+9x-5 }\]

OpenStudy (nicholausblackmon):

with you help me with mine before i help you with yours?

OpenStudy (nicholausblackmon):

because i already know the answer

OpenStudy (anonymous):

Is it a recent question you posted?

OpenStudy (nicholausblackmon):

im about to ask it

OpenStudy (anonymous):

I can't promise you I'll know how to solve, but I can try

OpenStudy (nicholausblackmon):

okay. Brb taking a screen shot

OpenStudy (nicholausblackmon):

its this one

OpenStudy (nicholausblackmon):

#13

OpenStudy (anonymous):

Oooooo, it's geometry...I SUCK at geometry.

OpenStudy (anonymous):

Never was a big fan of it, failed the class once, pass it the second time online.

OpenStudy (anonymous):

You still need help?

OpenStudy (anonymous):

Yes I do @manhani

OpenStudy (anonymous):

so you set the denominator to equal 0

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

Yes, that's as far as I knew to go.

OpenStudy (anonymous):

After that, I didn't know what to do

OpenStudy (anonymous):

So what happens when you set 2x^2+9x-5=0?

OpenStudy (anonymous):

Wouldn't you try to rid of the 2 in 2x^2 and get x^2 by itself?

OpenStudy (anonymous):

you dont have to, and you cant in the particular equation

OpenStudy (anonymous):

Because you cant divide 9 or 5 by 2

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So what times what will give you -10 and add up to 9

OpenStudy (anonymous):

I mean 10 and1

OpenStudy (anonymous):

-1*

OpenStudy (anonymous):

so its (x+5)(2x-1)=0

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

then you solve it from there and plug it in to the original equation to get a#/0 and if you get 0/0 it is not a vertical asym.

OpenStudy (anonymous):

So x=-5 and x=1/2

OpenStudy (anonymous):

yea and plug it in the original equation to get a #/0 and if you get 0/0 it is not a VA

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

Yes I do :)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

here

OpenStudy (anonymous):

https://www.youtube.com/watch?v=_qEOZNPce60

OpenStudy (anonymous):

Best of luck

OpenStudy (anonymous):

Thank you very much!! I truly do appreciate it! :)